# Stability-Aware Training of Neural Network Interatomic Potentials with Differentiable Boltzmann Estimators

### v1

@@ 2.1 Boltzmann Estimator
In some scenarios, unphysical configurations can occur within localized regions of the simulation domain, such as collisions between two molecules in a large condensed-phase system. Due to spatial averaging, global observables $g(\Gamma)$ may be insensitive to these localized events, limiting the ability to identify unphysical states. To address this, we propose the $N$-sample Localized Boltzmann Estimator. Here, the global energy $U_{\theta}(\Gamma)$ and observable $g(\Gamma)$ are replaced with local versions $U_{\theta}(\gamma)$ and $g(\gamma)$, where $\gamma$ denotes a local neighborhood of $n<N$ atoms. NNIPs compute global energies by summing individual atomic energies, and we can leverage this to compute $U_{\theta}(\gamma)$ by summing only over atoms in the local neighborhood. See Supplementary Section A.2 for a formal definition of the localized estimator. A concrete example of localized instability, along with the use of our localized Boltzmann estimator to correct it, will be presented in the context of an all-atom water system in Results, Section 3.4.
@@
The Boltzmann estimators defined earlier have several advantages over explicit differentiation through the MD simulation to train the NNIP. Due to the use of independent samples, the estimator avoids numerical instability associated with differentiating through long, continuous trajectories. Further, the most computationally expensive component of the estimator, the gradient of the potential energy $\nabla_{\theta}U_{\theta}(\Gamma)$, is independent of the observable $g(\Gamma)$. This means it can be reused, allowing efficient training to match multiple observables simultaneously. Finally, our localized formulation of the estimator enables finer control over isolated instabilities, providing a more informative gradient with which to optimize the NNIP.
@@ 2.2 Stability-Aware Boltzmann Estimator (StABlE) Training
Stability-Aware Boltzmann Estimator (StABlE) Training begins with conventional supervised pre-training of a NNIP on a dataset of structures from a high-fidelity, constant temperature MD simulation, paired with reference energy and forces values. The method then proceeds by alternating between two major phases, simulation and learning, as illustrated in Figure 2.
@@
The simulation phase aims to explore the molecular phase space and pinpoint regions where the NNIP becomes unstable. We sample $R$ equilibrium states from the training dataset as initial conditions for separate MD trajectories (replicas). Using the pre-trained NNIP $U_{\theta}$, we run MD simulations on these replicas in parallel for $t$ timesteps. We apply a predetermined stability criterion (see Supplementary Section A.5 for details) to each replica, freezing those marked unstable at their current states. Simulations continue for the remaining replicas, with stability checks every $t$ timesteps. Once a specified fraction of replicas become unstable, we rewind all unstable replicas by $t$ timesteps. This ensures that further simulation for $t$ timesteps will trigger instability, which is then corrected in the next phase.
@@
The learning phase aims to refine the NNIP to correct the previously encountered instabilities. Starting from the near-unstable configurations obtained in the simulation phase, we perform MD simulation for $t$ timesteps, sampling every $S^{th}$ state to obtain uncorrelated samples. Using the sampled states $\Gamma^{(1)},\ldots,\Gamma^{(N_{d})}$, where $N_{d}=\frac{t}{S}$, we compute the observable loss function (Equation 4). To update the NNIP parameters $\theta$, we compute an unbiased estimator of the loss gradient via the Boltzmann Estimator (Section 2.1) and use it to perform a single step of gradient descent. We then reset all replicas to their original near-unstable states, simulate with the updated NNIP, recompute the loss and gradient estimator with the newly sampled states, and again update the NNIP weights. This process is repeated until the fraction of unstable replicas drops below a predetermined threshold. When this occurs, the learning phase has concluded, and we begin a new simulation phase starting from the endpoints of the last learning phase. We continue alternating between simulation and learning phases until a predetermined computational budget is reached, at which point StABlE Training has concluded. See Supplementary Section A.3 for a formal algorithmic description of the StABlE Training procedure.
@@
In practice, the mapping between a sparse set of system observables and a potential energy function is non-unique [undefaw]. Consequently, learning with observables alone is underconstrained. To combat this, we regularize the observable loss function (Equation 4) with the energy and forces loss function (Equation 3). The final StABlE loss function thus becomes,
@@
where $\mathcal{L}_{\text{obs}}$ and $\mathcal{L}_{\text{QM}}$ were defined in Equations 4 and 3 respectively, and the hyperparameter $\lambda$ controls the strength of the regularization. Crucially, $\mathcal{L}_{\text{QM}}$ is still only computed over the static training dataset $\mathcal{D}_{\text{train}}$, and not over new structures explored during MD simulation. Therefore, the regularization introduces no additional quantum-mechanical supervision.
@@ 3 Results
[…]
We present the results of our StABlE training procedure on three different systems and NNIP architectures: 1) an aspirin molecule with a SchNet [undefaae] architecture (Sections 3.1 - 3.2), 2) an Ac-Ala3-NHMe tetrapeptide with a NequIP [undefd] architecture (Section 3.3), and 3) an all-atom water system with a GemNet-T [undefw] architecture (Section 3.4). For systems 1) and 2), we use the Boltzmann estimator presented in Equation LABEL:eq:estimator. For system 3), we employ our localized Boltzmann estimator (Equation 8) to correct isolated instabilities. We focus our evaluation of NNIPs on their ability to produce stable MD simulations which accurately recover key system observables. For all systems, stability thresholds are chosen to be very large relative to the scale of typical fluctuations in reference simulations at the chosen temperature. Simulations which cross the threshold thus cannot reasonably re-enter the manifold of stable configurations (see Supplementary Section A.5 for more details). We compute simulation observables by averaging over sampled states from the trajectory corresponding to the median stability improvement (over all replicas) between the baseline and StABlE-trained NNIP. More detailed information about NNIP architectures, training, simulation, and evaluation can be found in Supplementary Sections A.6, A.7, and A.8.
@@ 3.1 Aspirin Molecule
Aspirin (chemical formula $C_{9}H_{8}O_{4}$) is the largest molecule from MD17 [undefq], a widely used benchmark dataset for atomistic simulations which contains energy and forces calculations computed at the PBE+vdW-TS [undefaam] level of theory. Consisting of 21 atoms, aspirin has been shown to be the most challenging molecule in MD17 for state-of-the-art NNIPs to simulate stably [undefv].
Problem setup.
@@
We pre-train a SchNet [undefaae] model on the energy and forces matching objective (Equation 3) using a subset of 1,000 aspirin structures from the reference dataset. After convergence, we begin StABlE Training with the global Boltzmann estimator, using 128 parallel replicas at a temperature of 500K. By simulating hundreds of replicas in parallel, we expose the NNIP to a comprehensive range of failure modes much more quickly than possible with a single replica. Following [undefv], we use a maximum bond length deviation criterion to detect instability in the simulations (see Supplementary Section A.5 for details). We use the distribution of interatomic distances, $h(r)$, as our training observable, with the $h(r)$ computed over structures in the training dataset serving as the reference (see Supplementary Section A.4 for details). This choice is motivated by the observation that unphysical bond stretches constitute the majority of failure cases in aspirin simulations. The $h(r)$ is sensitive to these abnormalities, and is thus an informative optimization criterion. We perform four complete StABlE cycles of simulation and learning with a SGD optimizer.
Results.
We evaluate our final models by performing constant-temperature MD simulations, starting from 256 initial structures not seen during training. Figure 3a shows the progression of StABlE Training. During the simulation phase, the learning rate is zero and the fraction of unstable replicas steadily increases. When the upper instability threshold is reached, learning commences. The fraction of unstable replicas steadily decreases, confirming that optimization of the interatomic distance distribution produces the desired stability improvement. A decaying learning rate schedule is employed to avoid overfitting to the current batch of replicas.
A SchNet model trained via our stability-aware approach is significantly more stable during MD simulation than models trained only on the conventional QM energy and forces objective function (Figure 3b). As a result of our training procedure, the mean stable simulation time increases from 35 to 140 picoseconds. Even when StABlE training is employed after training on only on 1,000 reference structures, it is more stable than models trained conventionally on 10,000 and 50,000 reference structures, highlighting the effectiveness of our method in improving stability without reliance on any additional QM reference data. We also observe that simply reducing the simulation timestep does not reliably lead to an improvement in the stability of simulations produced by conventionally trained models (see Supplementary Section A.11 for details).
@@ Results.
Simulations from our model closely recover the true distribution of interatomic distances, while the distributions produced by models solely trained on QM data are noisier, due to the limited stable simulation time (Figure 3c). We also test the ability of our StABlE-trained model to recover the velocity autocorrelation function (VACF), a fundamental dynamical observable not seen during training (Figure 33). We see that the quality of the recovered VACF remains similar after StABlE Training, with lower variance over the simulated replicas (see Supplementary Section A.12 for a sample VACF computed from the simulations). The preservation of a held-out dynamical observable suggests that StABlE Training does not achieve stability improvements by simply restricting the model around a narrow regime corresponding to the reference $h(r)$. If this were the case, characteristic dynamic fluctuations around the reference $h(r)$ would be suppressed, leading to an inaccurate VACF. Figure 3e depicts aspirin structures sampled from simulations during the course of a single learning phase of StABlE Training, demonstrating that unphysical bond stretching in these structures is resolved by the training procedure.
@@ 3.2 Temperature generalization
Figure 4: Testing temperature generalization for aspirin. We test the generalization of StABlE Training by performing 256 parallel MD simulations at 350 Kelvin and 700 Kelvin with a SchNet model which was StABlE-trained at 500 Kelvin on aspirin from MD17. (a, b) Fraction of unstable replicas as a function of simulation time for a SchNet NNIP trained on 1,000 mechanical (QM) energy and forces reference data points (blue). Applying StABlE Training to this model yields significantly more stable simulations (orange). At 350K, before StABlE Training, $70\%$ of replicas are unstable after 300 picoseconds of simulation. After StABlE Training, <<5% of replicas are unstable. (c, d) Distribution of interatomic distances ($h(r)$) from an ab-initio molecular dynamics (AIMD) simulation (black), a SchNet model trained on 1,000 QM data points (blue), and a SchNet model trained on 1,000 QM datapoints, followed by our StABlE Training procedure (orange). The StABlE-trained SchNet model more accurately recovers the true $h(r)$ at both 350K and 700K. At 350K, the baseline model trained only on QM data incorrectly estimates several peaks in the $h(r)$, while at 700K the baseline model produces a noisy $h(r)$ due to its limited stable simulation time (inset shows difference between the predicted and reference $h(r)$). The AIMD reference $h(r)$ is computed by Boltzmann-reweighting of independent samples from the 500K training dataset.
Problem setup.
We explore our method’s generalization to different thermodynamic conditions. To do so, we perform 256 parallel MD simulations at 350K and 700K using the SchNet model which was trained with our StABlE Training procedure at 500K. We use the same criterion based on bond length deviation to detect instability as in Section 3.1. We estimate the reference $h(r)$ at 350K and 700K by Boltzmann-reweighting samples from the original 500K training dataset (see Supplementary Section A.9 for details).
Results.
Aspirin simulations produced by a StABlE-Trained SchNet model are significantly more stable at both temperatures than baseline models trained only on QM reference data (Figure 4a, b). At 350K, the difference is particularly large. After 300 picoseconds of simulation, less than 5% of replicas are unstable in the StABlE-trained NNIP simulation, while over 75% replicas are unstable in the baseline NNIP simulation. The StABlE-trained model also recovers the true distribution of interatomic distances more closely than the baseline NNIP at both temperatures (Figure 4c, d). At 350K, the baseline NNIP is stable enough produce a smooth $h(r)$, but the structure is inaccurate, indicating incorrect sampling of the phase space. At 700K, the baseline NNIP produces a $h(r)$ which is close to the AIMD-produced distribution, but is noisy due to insufficient stable simulation time. This underscores the necessity for a NNIP to be both stable and accurate in order to be practically useful in MD simulation. StABlE Training improves both of these metrics for this system. As generalization to new thermodynamic conditions is of crucial importance for NNIPs to be useful in practical applications [undefah], we see the temperature-transferability of StABlE Training as an important strength of our method.

### v2

@@ 3.1 Boltzmann Estimator
In some scenarios, unphysical configurations can occur within localized regions of the simulation domain, such as collisions between two molecules in a large condensed-phase system. Due to spatial averaging, global observables $g(\Gamma)$ may be insensitive to these localized events, limiting the ability to identify unphysical states. To address this, we propose the $N$-sample Localized Boltzmann Estimator. Here, the global energy $U_{\theta}(\Gamma)$ and observable $g(\Gamma)$ are replaced with local versions $U_{\theta}(\gamma)$ and $g(\gamma)$, where $\gamma$ denotes a local neighborhood of $n<N$ atoms. Formally, define a neighborhood of $n$ atoms $\mathcal{N}=\{x_{1},x_{2},\ldots,x_{n}\mid x_{i}\in\mathbb{Z},1\leq x_{i}\leq N\ \text{for all}\ i=1,2,\ldots,n\}$. The local state $\gamma$ is defined as $\gamma=\{[r^{(\mathcal{N}_{1})};\ldots;r^{(\mathcal{N}_{n})}],[p^{(\mathcal{N}_{1})};\ldots;p^{(\mathcal{N}_{n}})]\}$. The local energy is then defined as $U_{\theta}(\gamma)=\sum_{i=1}^{n}U_{\theta}(\gamma^{(i)})$, where $\gamma^{(i)}=\{r^{(\mathcal{N}_{i})},p^{(\mathcal{N}_{i}})\}$ contains the position and momenta of the $i^{th}$ atom in the local neighborhood. The local energy is easily obtained by noting that MLFFs parameterize their global energy prediction $U_{\theta}$ as a sum over individual atomic energies. The localized estimator is thus given as follows:
Definition 2 ($N$-sample localized Boltzmann estimator).
Given $N$ i.i.d. samples of local states $\gamma_{1},\ldots,\gamma_{N}$, where each $\gamma_{i}$ is extracted from a global state $\Gamma_{i}\sim P_{\theta}(\Gamma)$, we define the $N$-sample localized estimator of the Jacobian $\frac{\partial\mathbb{E}_{\gamma\sim P_{\theta}(\gamma)}[g(\gamma)]}{\partial\theta}$ as $\mathcal{E}(\gamma_{1},\ldots,\gamma_{N})\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{N}{k_{B}T(N-1)}\left[\hat{\mathbb{E}}\left[g(\gamma)\right]\hat{\mathbb{E}}\left[{\nabla_{\theta}U_{\theta}(\gamma)}\right]^{\top}-\hat{\mathbb{E}}\left[{g(\gamma)\cdot\nabla_{\theta}U_{\theta}(\gamma)}^{\top}\right]\right]$ (5)
The localized estimator follows from the original Boltzmann estimator due to the fact that as a subset of the larger Boltzmann-distributed system, any local neighborhood also obeys a Boltzmann-distribution. In practice, we extract multiple local neighborhoods from each global state to increase the sample size and state space coverage (Section A.6). A concrete example of localized instability, along with the use of our localized Boltzmann estimator to correct it, will be presented in the context of an all-atom water system in Results, Section 4.4.
@@
Unlike active learning approaches, no additional quantum mechanical energy and forces calculations are required to compute our Boltzmann Estimators (all energy terms are model predictions). As a result, learning with the Boltzmann estimator is computationally efficient. Due to the use of independent samples, the estimator also avoids numerical instability and memory demands associated with differentiating through long, continuous trajectories. Further, the most computationally expensive component of the estimator, the gradient of the potential energy $\nabla_{\theta}U_{\theta}(\Gamma)$, is independent of the observable $g(\Gamma)$. This means it can be reused, allowing efficient training to match multiple observables simultaneously.
@@ 3.2 Stability-Aware Boltzmann Estimator (StABlE) Training
Stability-Aware Boltzmann Estimator (StABlE) Training begins with conventional supervised pre-training of a MLFF on a reference dataset of energy and forces, using the loss defined in Equation 1. The method then proceeds by alternating between two major phases, simulation and learning, as illustrated in Figure 2.
@@
The simulation phase aims to explore the molecular phase space and pinpoint regions where the MLFF becomes unstable. We sample $R$ equilibrium states from the training dataset as initial conditions for separate MD trajectories (replicas). Using the pre-trained MLFF $U_{\theta}$, we run MD simulations on these replicas in parallel for $t$ timesteps. Due to our use of efficient vectorized GPU operations, simulating many replicas in parallel has similar computational cost to simulating a single replica. We apply a predetermined stability criterion (Supplementary Section A.5) to each replica, freezing those marked unstable at their current states. Simulations continue for the remaining replicas, with stability checks every $t$ timesteps. Once a specified fraction of replicas become unstable, we rewind all unstable replicas by $t$ timesteps. This ensures that further simulation for $t$ timesteps will trigger instability, which is then corrected in the next phase. We note that many other strategies exist to explore phase space and find unstable regions, including Diffusion Monte Carlo with fictitious masses Li et al. (2021) or uncertainty-based sampling Kulichenko et al. (2023). StABlE Training can be flexibly used with any of these strategies without impacting the learning stage (described next) of the procedure.
@@
The learning phase aims to refine the MLFF to correct the previously encountered instabilities. Starting from the near-unstable configurations obtained in the simulation phase, we perform MD simulation for $t$ timesteps, sampling every $S^{th}$ state to obtain uncorrelated samples. Using the sampled states $\Gamma_{1},\ldots,\Gamma_{N_{d}}$, where $N_{d}=\frac{tR}{S}$, we compute the observable loss function (Equation 2). To update the MLFF parameters $\theta$, we compute an unbiased estimator of the loss gradient via the Boltzmann Estimator (Section 3.1) and use it to perform a single step of gradient descent. We then reset all replicas to their original near-unstable states, simulate with the updated MLFF, recompute the loss and gradient estimator with the newly sampled states, and again update the MLFF weights. This process is repeated until the fraction of unstable replicas drops below a predetermined threshold. When this occurs, the learning phase has concluded, and we begin a new simulation phase starting from the endpoints of the last learning phase. We continue alternating between simulation and learning phases until a predetermined computational budget is reached, at which point StABlE Training has concluded. See Supplementary Section A.3 for a formal algorithmic description of the StABlE Training procedure.
@@
In practice, the mapping between a sparse set of system observables and a potential energy function is non-unique Noid (2013). Consequently, learning with observables alone is underconstrained. To combat this, we regularize the observable loss function (Equation 2) with the energy and forces loss function (Equation 1). The final StABlE loss function thus becomes
@@
where $\mathcal{L}_{\text{obs}}$ and $\mathcal{L}_{\text{QM}}$ were defined in Equations 2 and 1 respectively, and the hyperparameter $\lambda$ controls the strength of the regularization. Crucially, $\mathcal{L}_{\text{QM}}$ is still only computed over the original training dataset $\mathcal{D}_{train}$, and not over new structures explored during MD simulation. Therefore, the regularization requires no additional quantum mechanical energy and forces calculations.
@@ 4 Results
We present the results of StABlE Training on the aspirin molecule with a SchNet MLFF (Section 4.1), the Ac-Ala3-NHMe tetrapeptide with a NequIP MLFF (Section 4.3), and an all-atom water system with a GemNet-T MLFF (Section 4.4).
@@ 4.1 Aspirin Molecule
Figure 3: Aspirin simulation with StABlE Training. (a) Alternation of simulation and learning phases during StABlE Training with 128 parallel replicas. The simulation phases correspond to regions where the fraction of unstable replicas increases, while the learning phases correspond to regions where learning occurs and the fraction of unstable replicas decreases. (b) Stable simulation time of 256 parallel aspirin trajectories from SchNet MLFFs. Applying StABlE Training yields significantly more stable simulations than models trained only on energies and forces, surpassing models trained on 50$\times$ more QM data. (c) Distribution of interatomic distances ($h(r)$) from MLFF simulations. A StABlE-trained SchNet model closely recovers the true distribution of interatomic distances, while the model trained only on QM reference data produces a noisier $h(r)$ because it cannot stably simulate the system for longer time periods. Inset shows difference between predicted and reference $h(r)$. (d) Distribution of velocity autocorrelation function (VACF) mean absolute error (MAE). StABlE Training yields a reduction in variance across replicas. (e) Aspirin structures sampled over epochs of a single learning phase of StABlE Training. There is a clear progression as unstable configurations become stable.
@@
Aspirin (chemical formula $C_{9}H_{8}O_{4}$) is the largest molecule from MD17 Chmiela et al. (2017), a widely used benchmark dataset for atomistic simulations which contains energy and forces calculations computed at the PBE+vdW-TS Tkatchenko & Scheffler (2009) level of theory. Consisting of 21 atoms, aspirin has been shown to be the most challenging molecule in MD17 for state-of-the-art MLFFs to simulate stably Fu et al. (2022). We pre-train a SchNet Schütt et al. (2018) model on the energy and forces matching objective (Equation 1) using a subset of 1,000 aspirin structures from the reference dataset. After convergence, we begin StABlE Training with the global Boltzmann estimator, using 128 parallel replicas at a temperature of 500K. By simulating hundreds of replicas in parallel, we expose the MLFF to a comprehensive range of failure modes much more quickly than possible with a single replica. Following Fu et al. (2022), we use a maximum bond length deviation criterion to detect instability in the simulations (Section A.5). We use the distribution of interatomic distances, $h(r)$, as our training observable, with the $h(r)$ computed over structures in the training dataset serving as the reference (Section A.4). This choice is motivated by the observation that unphysical bond stretches constitute the majority of failure cases in aspirin simulations. The $h(r)$ is sensitive to these abnormalities, and is thus an informative optimization criterion. We perform four complete StABlE cycles of simulation and learning with a SGD optimizer. We evaluate our final models by performing 1 nanosecond constant-temperature MD simulations, starting from 256 initial structures not seen during training. Figure 3a shows the progression of StABlE Training. During the simulation phase, the learning rate is zero and the fraction of unstable replicas steadily increases. When the upper instability threshold is reached, learning commences. The fraction of unstable replicas steadily decreases, confirming that optimization of the interatomic distance distribution produces the desired stability improvement.
Figure 4: Testing temperature generalization for aspirin. (a, b) Fraction of unstable replicas as a function of simulation time for SchNet MLFFs at 350K and 700K. Applying StABlE Training yields significantly more stable simulations than a baseline SchNet model. (c, d) Distribution of interatomic distances ($h(r)$) from SchNet MLFFs. The StABlE-trained SchNet model more accurately recovers the true $h(r)$ at both 350K and 700K (the inset shows difference between the predicted and reference $h(r)$). The AIMD reference $h(r)$ is computed by Boltzmann-reweighting of independent samples from the 500K training dataset.
@@
A SchNet model trained via our stability-aware approach is significantly more stable during MD simulation than models trained only on the conventional QM energy and forces objective function (Figure 3b). As a result of our training procedure, the median stable simulation time increases from 42 to 602 picoseconds. Even when StABlE training is employed after training on only on 1,000 reference structures, it is more stable than models trained conventionally on 10,000 and 50,000 reference structures, highlighting the effectiveness of our method in improving stability without reliance on any additional QM reference data. We also observe in Supplementary Section A.12 that simply reducing the simulation timestep by a factor of ten does not eliminate instability in simulations produced by conventionally trained models, and reducing instability further may require impractically small timesteps. The stability gains from reducing the timestep are smaller than those resulting from StABlE Training, suggesting that our approach could enable taking larger timesteps and accelerate simulations. Simulations from our model closely recover the true distribution of interatomic distances, while the distributions produced by models solely trained on QM data are noisier, due to the limited stable simulation time (Figure 3c). We also test the ability of our StABlE-trained model to recover the velocity autocorrelation function (VACF), a fundamental dynamical observable not seen during training (Figure 3d). We see that the quality of the recovered VACF remains similar after StABlE Training, with lower variance over the simulated replicas (see Supplementary Section A.13 for a sample VACF computed from the simulations). The preservation of a held-out dynamical observable suggests that StABlE Training does not achieve stability improvements by simply restricting the model around a narrow regime corresponding to the reference $h(r)$. If this were the case, characteristic dynamic fluctuations around the reference $h(r)$ would be suppressed, leading to an inaccurate VACF. Figure 3e depicts aspirin structures sampled from simulations during the course of a single learning phase of StABlE Training, demonstrating that unphysical bond stretching in these structures is resolved by the training procedure.
@@ 4.2 Temperature generalization
We explore our method’s generalization to different thermodynamic conditions. To do so, we perform 256 parallel MD simulations at 350K and 700K using the SchNet model which was trained with our StABlE Training procedure at 500K. We use the same criterion based on bond length deviation to detect instability as in Section 4.1. We estimate the reference $h(r)$ at 350K and 700K by Boltzmann-reweighting samples from the original 500K training dataset (Section A.9). Aspirin simulations produced by a StABlE-Trained SchNet model are significantly more stable at both temperatures than baseline models trained only on QM reference data (Figure 4a, b). At 350K, the difference is particularly large. After 1 nanosecond of simulation, virtually no replicas are unstable in the StABlE-trained MLFF simulation, while approximately 80% of replicas are unstable in the baseline MLFF simulation. The StABlE-trained model also recovers the true distribution of interatomic distances more closely than the baseline MLFF at both temperatures (Figure 4c, d). At 350K, the baseline MLFF is stable enough produce a smooth $h(r)$, but the structure is inaccurate, indicating incorrect sampling of the phase space. At 700K, the baseline MLFF produces a $h(r)$ which is close to the AIMD-produced distribution, but is noisy due to insufficient stable simulation time. This underscores the necessity for a MLFF to be both stable and accurate in order to be practically useful in MD simulation. StABlE Training improves both of these metrics for this system. As generalization to new thermodynamic conditions is of crucial importance for MLFFs to be useful in practical applications Kovács et al. (2021), we see the temperature-transferability of StABlE Training as an important strength of our method.